OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -9) and a directrix y = 9.

OpenStudy (anonymous):

Answer Choices: A. y = -1/9x^2 B. y^2 = -36x C. y = -1/36x^2 D. y^2 = -9x

OpenStudy (anonymous):

What do you think the answer is? Do you know what a focus is?

OpenStudy (serenity74):

What's the formula for parabola? How can you plug it into the formula?

OpenStudy (anonymous):

The formula is (x−h)^2=4p(y−k)

OpenStudy (anonymous):

@tom982 Where (h,k) is the vertex

OpenStudy (anonymous):

and focus is either (h, k+p) or (h+p ,k)

OpenStudy (anonymous):

That's all I know

OpenStudy (anonymous):

Anybody?

OpenStudy (whpalmer4):

Let's work through it with the material you provided. \[ (x−h)^2=4p(y−k)\] focus at \( (0, -9)\) and directrix of \(y = 9\) Any idea where the vertex will be?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

I don't know how to get the vertex from that information.

OpenStudy (whpalmer4):



OpenStudy (whpalmer4):

yo, @turmoilx2 are you going to take part here?

OpenStudy (anonymous):

hmmmmmmmmm.....

OpenStudy (anonymous):

Yes. I am replying. @whpalmer4

OpenStudy (anonymous):

I just don't see what I'm supposed to do with the information yet.

OpenStudy (whpalmer4):

Do you know the definition of what a parabola is? It is a set of points that satisfy some rule, do you know what the rule is?

OpenStudy (whpalmer4):

sorry, open study is not behaving well for me, going to try restarting my browser, brb...

OpenStudy (mathmale):

I'd suggest that you attempt to graph this parabola. Draw a set of axes and locate the given point and the given line. Identify the point, the line and the vertex of your parabola. Even a quick sketch would be helpful. Your job will be to relate distances in this drawing to the equation of a parabola. What does "p" in your formula represent? If you're not sure, look it up, please. Try the Internet and search for "equations of a parabola." Your graph should give you the value of p by inspection.

OpenStudy (anonymous):

OS is not behaving well for me either. That is why I haven't been participating in the solving of my question.

OpenStudy (mathmale):

Important: The vertex of a parabola lies exactly halfway between the directrix (a line) and the focus (a point). If the directrix is the horiz. line y=9 and the focus is at (0,-9), what are the coordinates of the vertex? Again, I ask you to graph this situation. Show the directrix, the focus and the vertex.

OpenStudy (anonymous):

 Here's my sketch

OpenStudy (mathmale):

Please see the following on the 'Net: https://www.google.com/search?sourceid=chrome-psyapi2&ion=1&espv=2&es_th=1&ie=UTF-8&q=directrix%20parabola&oq=directrix%20parabola&aqs=chrome..69i57j0l5.3199j0j7 The very first illustration shows that the vertex is halfway between the directrix and the focus. p is the distance between the focus and the vertex. What is p here?

OpenStudy (mathmale):

Will YOUR parabola open up or down? You are given the directrix: y=9, and the focus: (0,-9). The vertex is in between. UP or DOWN?